# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-modal Consistency in Visual-Stimulated Narratives

### v1

@@ IV Approach
In this work, we propose a comprehensive analysis for the automatic detection of Neurocognitive Disorders based on visual-stimulated narratives, encompassing both microstructural and macrostructural analysis. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features selected from previous literature. On the other hand, macrostructural analysis, presented in Sections IV-B to IV-D, introduces three distinct methodologies to capture global and temporal patterns with varying focal points.
@@
A1 # Total pause # pauses
@@
A6 Pause occurrence rate # pauses / # characters
@@
A9 Articulation rate # characters spoken in a second
@@
L1 % Stop words # stop words / # words
L2 % Filler pauses # filler words / # words
L3 % Lexical filler # lexical filler / # words
L4 % Repetition # repeated words / # words
L5 % Pronoun verb change # pronoun to verb changes / # words
L6 % Pronoun noun change # pronoun to noun changes / # words
L7 % Pronoun # pronouns / # words
L8 % Verb # verb / # words
@@
L10 % Adj # adjectives / # words
L11 % Pronoun verb # pronouns / # verbs
L12 % Pronoun noun # pronouns / # noun
L13 % Type of token (TTR) # different tokens / # tokens
IV-A Traditional Acoustic and Linguistic Analysis
Based on the literature review, the microstructural features have been categorized into two main groups: acoustic features and linguistic features. We select 9 acoustic features and 13 linguistic features from the literature that are suitable for Cantonese and CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all audios with FullSubNet+ [69] denoiser and peak-normalization to ensure internal consistency and avoid Clever Hans Effect [70].
The rows of A1-A9 in Table II show the name and the description of the acoustic features, related to pause (A1-A6), and speech duration (A7-A8) and speaking rate (A9). To extract these features, we first employ Voice Activity Detection (VAD) using the Feedforward Sequential Memory Network, as described in the paper [71], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected by the silence gap between two adjacent voiced segments when the duration exceeds the duration of a syllable for this speaker, which is calculated by dividing the total voiced duration by the total number of characters for each Cantonese participant.
The linguistic features are extracted from the automatic transcriptions of participants obtained using the ASR model SenseVoice-Small [72]. The rows of L1-L13 in Table II show the name and the corresponding description of each feature. To extract these linguistic features from Cantonese transcripts, we utilized PyCantonese [^1] https://pycantonese.org to do word segmentation and POS auto-tagging, combining vocabulary from Jieba [^2] https://github.com/fxsjy/jieba.
@@ IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a subject’s narrative. To this end, we identify keywords associated with the task and measure the coverage in the subject’s narratives. Specifically, we leverage the TF-IDF approach from information retrieval to compute the relative importance of each word within HC groups. After selecting the top 100 words based on TF-IDF weights and excluding punctuation and stopwords, these words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story and the Narrative Scoring Scheme (NSS) proposed by [24]. The groups include introduction setting words (e.g., one day, morning), character words (e.g., girl, dog, rabbit), object words (e.g., park, home, stone), and action words (e.g., play, find, walk). Coverage (abbreviated as “cvg”) for each category is then computed as features for analysis. While these coverage metrics effectively capture whether local narrative elements are addressed, they lack insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1,2,3,4), METEOR, and ROUGE-L, are employed to quantify the degree of alignment between subject narratives and reference descriptions. For this purpose, we engaged six colleagues to craft curated reference descriptions of the Rabbit Story that align with the rubric of NSS, in the aspects of introduction, character development, mental and emotion states, referencing, event&reaction, cohesion, and conclusion. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each subject, providing additional features that reflect global narrative coherence. With these reference-based features, we can perform macrostructural analysis that evaluates the coherence against some gold standards; however, these features do not account for temporal changes in topics or the cohesion of narrative over time, which is another important part of microstructural analysis.
@@ IV-C DTM-based Temporal Analysis
To understand how DTM captures these temporal dynamics, it is important to first grasp the basics of topic modeling. Topic model provides a probabilistic framework to undercover the latent semantic structure within textural data. While the arbitrarily generated text may randomly be distributed across documents, the topic model assumes that words in natural languages are distributed around meaningful semantic topics. Specifically, document generation is assumed to be based on a probabilistic combination of underlying topics, where each topic is represented by a probabilistic distribution over vocabulary words. In the classical Latent Dirichlet Allocation (LDA), these distributions are assumed to be static and drawn from Dirichlet priors.
Dynamic Topic Models (DTM) is an extension of classical topic models, it introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing time dependency $t$ in topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
@@
In order to apply the DTM model to narrative analysis for NCD detection, we evenly split each participant’s ASR transcription into 15 slices, assuming that each image leads to approximately similar words of spoken narratives. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
@@
Fig. 3 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 3(a), include “house” and “garden” in topic “A”, “ant” and “tree” in topic “B”, “goose” and “caw” in topic “C”, “pond” in topic “D”, and “many” “rabbits” in topic “E”. These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 3(b), are more chaotic, e.g., topic “A” and “B” are mostly overlapped with “rabbit” and “dog”, while “stable” in “C” and “fall” in “E” are both derived from images 10-11.
@@
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC individuals mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This implies that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the story to a satisfying closure. For the dynamic proportions of these topics over time, i.e., the Gaussian Process, it is observable that the HC group exhibits more distinct peaks and smoother transitions between topics over time, while the NCD group shows more erratic changes with less pronounced peaks.
@@
Based on these findings, we proposed several DTM-related statistics for NCD detection. The first metric, termed “DTM-based topic consistency”, is derived from the hit rate of the top-10 topic words each time, which are derived from a DTM pre-trained on all narratives. This statistic measures how consistently the subject’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “DTM-based cyclical pattern”, which is simply a boolean value indicating whether the word “home” appears in the latter half of each narrative.
We also proposed several metrics by calculating the dynamic properties of topic evolution, including standard deviation of the proportions over time, peak-to-peak range (distance between positions of the highest and lowest points), and rate of change (averaged Euclidean distance of adjacent topic distributions over time).
@@ IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and visual-stimulated narratives. This limitation impairs topic representation model, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [42]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [76, 77], we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 4.
@@
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising text-tower and visual-tower to process different modalities. In this work, we adopted the Chinese-CLIP [78] model, which consists of visual-tower ViT and text-tower RoBERTa. Chinese-CLIP is pre-trained in a contrastive learning paradigm that aims to match the images with their corresponding descriptions in a multi-modal embedding space, specifically tailored for the Chinese language (including Cantonese). We adopt this model because it can effectively align the embeddings of images and Cantonese text in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the respective embeddings $E_{I}$ and $E_{T}$ as follows:
@@
where $H$ is the hidden dimension of the extracted embeddings, $K$ is the temporal length of the investigated text sequence, and $J$ is the number of image frames, e.g., $J=1$ for the ADReSS dataset and $J=15$ for the CU-MARVEL-RABBIT corpus.
Figure 4: (a) The proposed model TITAN. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN. TITAN takes images and text narratives as inputs to predict NCDs, with Attention using RoPE to focus on useful positions.
@@
$E_{attn}=\text{softmax}\left(\frac{\text{Mask}(Q^{\prime}K^{\prime T})}{\sqrt{H^{\prime}}}\right)V$
@@
In visual-stimulated narrative tasks such as Rabbit Story, narrative transcriptions must align with reference images. Participants with NCDs may misidentify animal names (anomia) or miss detailed scenarios (visual perception) in these images. For example, most NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The attention module in proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
TABLE III: Classification and Regression Results on CU-MARVEL-RABBIT. The quantity or dimension of the features is shown in parentheses. The star ‘*’ in ‘Sys.’ signifies the baseline systems. $\downarrow$ indicates lower is better. “Emb.” denotes embedding(s).
@@
1* Acoustics (9) SVM 0.5778 0.8029 0.4333 0.8667 0.6200 0.178 0.3206
2* Linguistics (13) SVM 0.2933 0.7296 0.1833 0.7333 0.4700 -0.0377 0.3602
3* Micro (22) SVM 0.5843 0.8204 0.4333 0.8966 0.6300 0.1904 0.3181
4 Reference-based (16) SVM 0.6735 0.7350 0.5500 0.8684 0.6800 0.2381 0.3086
5 DTM-based (5) SVM 0.6737 0.7792 0.5333 0.9143 0.6900 0.2508 0.3060
6 Macro (21) SVM 0.6804 0.7821 0.5500 0.8919 0.6900 0.2864 0.2987
7 All statistics (43) SVM 0.6947 0.7942 0.5500 0.9429 0.7100 0.2570 0.3048
@@
where $C$ is the prediction dimension (number of classes for classification tasks and 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the fully connected layer’s weight and bias, and $\sigma$ is an activation function (softmax for classification tasks and sigmoid for the regression task).
@@
The TITAN model, with its robust mechanisms for feature extraction, embedding fusion, and Attention-based contextual alignment, effectively integrates text and image information to evaluate the consistency and temporal flow in narratives. As shown in Fig. 4(c), the heatmap illustrates how the model captures the alignment between text chunks (x-axis) and a sequence of images (y-axis). The varying intensity of the colors represents the degree of alignment between the image and textual content, and the diagonal intensity represents the temporal consistency and flow of the narrative. Bright spots indicate high alignment, showing that text chunks and images correlate strongly at those points, while the diagonal progression reflects the sequence in which events or scenes unfold coherently over time. Moreover, compared to Fig. 3(c, d) using term-based topic representation, Fig. 4(c) demonstrates a more fine-grained temporal alignment between narratives and topics/images, which highlights the efficiency of representing topics with images.

### v2

@@ IV Approach
In this work, we propose a comprehensive analysis for the automatic detection of Neurocognitive Disorders based on VSNs, encompassing both microstructural and macrostructural analysis. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features selected from previous literature. On the other hand, macrostructural analysis, presented in Sections IV-B to IV-D, introduces three distinct methodologies to capture global and temporal patterns with varying focal points.
IV-A Traditional Acoustic and Linguistic Analysis
Based on the literature review, the microstructural features have been categorized into two main groups: acoustic features and linguistic features. We select 10 acoustic features and 13 linguistic features from the literature that are basic, universal and suitable for Cantonese in the CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all audios with FullSubNet+ [72] denoiser and peak-normalization to ensure internal consistency and avoid the Clever Hans Effect (where the model exploits irrelevant or unintended shortcuts like noise) [73].
The rows of A1-A10 in Table II list the acoustic features, related to pause (A1-A6), speech duration (A7-A8), articulation rate (A9) and speaking rate (A10). The articulation rate measures phonetic production speed as the number of syllables per voiced duration (excluding pauses), while the speaking rate quantifies overall delivery speed as the number of syllables per total duration (including pauses), thereby reflecting real-time communication fluency [74].
To extract these features, we first employ Voice Activity Detection (VAD) using the pretrained Feedforward Sequential Memory Network, as described in the paper [75], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected by the silence gap between two adjacent voiced segments when the duration exceeds the duration of a syllable for this speaker, which is calculated by dividing the total voiced duration by the total number of syllables for each Cantonese participant.
The linguistic features are extracted from the automatic transcriptions of participants obtained using the pretrained Automatic Speech Recognition (ASR) model SenseVoice-Small [76], which demonstrates a Word Error Rate (WER) of 31.37% on the ADReSS test set and a Character Error Rate (CER) of 34.22% on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus. The rows of L1-L13 in Table II show the name and the corresponding description of each feature. To extract these linguistic features from Cantonese transcripts, we utilized pretrained BERT base models for word segmentation [^3] https://huggingface.co/ckiplab/bert-base-chinese-ws and POS auto-tagging [^4] https://huggingface.co/ckiplab/bert-base-chinese-pos, with F1 scores of 82.69% and 85.11% respectively on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus.
@@ IV-A Traditional Acoustic and Linguistic Analysis
A1 # pauses The number of pauses
@@
A6 Pause occurrence rate # pauses / # syllables
@@
A9 Articulation rate # syllables / voiced dur
A10 Speaking rate # syllables / total dur (include pauses)
@@
L1 # words The number of words
L2 % Stop words # stop words / # words
L3 % Filled pauses # filler words / # words
L4 % Lexical filler # lexical filler / # words
L5 % Backchannel # backchannel words / # words
L6 % Repetition # repeated words / # words
L7 % Adj # adjectives / # words
L8 % Adv # adverbs / # words
@@
L10 % Pronoun # pronouns / # words
L11 % Verb # verb / # words
L12 % Func # functional words / # words
L13 % CTTR # types / rooted(2 * # tokens)
@@ IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a subject’s narrative. To this end, we identify keywords associated with the task and measure the coverage in the subject’s narratives. Specifically, we rank words from the HCs group (excluding punctuation and stopwords) based on their relevance to visual stimuli. This relevance is quantified through cosine similarity between global image embeddings (from the global special token [CLS]) and text embeddings (via mean pooling), both extracted from the final projection layers of the visual and textual encoders in the Chinese-CLIP model [77], which is pretrained via a contrastive learning paradigm that aligns images with corresponding textual descriptions on diverse Chinese variations (including the Cantonese dialect). As the textual and visual modalities are projected into a shared multimodal embedding space, the Chinese-CLIP model inherently enables cross-modal similarity measurement. These derived top-ranked words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story, and the Narrative Scoring Scheme (NSS) proposed by [24], including general settings or introduction (e.g., daytime/one day, garden/house/forest), characters or roles (e.g., girl, dog, rabbit), objects (e.g., sofa, box, stone), and actions (e.g., play, find, walk). Coverage (abbreviated as “cvg”, ranging from 0 to 1) for each category is then computed as features for analysis, focusing on the narrative content instead of linguistic form like PoS ratios. For example, the Rabbit Story features 7 characters, but the NCD narrative in Fig. 2(b) only mentions 4 of them, resulting in a “cvg_role” value of 4/7.
While these coverage metrics effectively capture whether local narrative elements are addressed, they lack insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1,2,3,4), METEOR, and ROUGE-L, are employed to quantify the degree of alignment between subject narratives and reference descriptions. For this purpose, we engaged six research team members, all thoroughly familiar with the NSS rubric [^5] The NSS rubric evaluates narratives across seven dimensions: introduction, character development, mental and emotional states, referencing, conflict/resolution (or event/reaction), cohesion, and conclusion. Each dimension uses a 0-5 scale, with concrete elaborations and examples provided for the “Proficient (5)”, “Emerging (3)”, and “Minimal/Immature (1)” levels., to craft reference descriptions of the Rabbit Story that exemplify the proficient criteria. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each subject, providing additional features that reflect global narrative coherence.
In short, we derive two types of features for reference-based analysis: (1) 10 domain-specific coverage metrics, including introduction, character, object, and action words, as well as high-frequency visual-related terms (nouns, verbs, pronouns, adjectives, adverbs, and all types of POS); and (2) 6 expert-aligned narrative similarity scores (BLEU-n, METEOR, ROUGE-L). These reference-based features enable macrostructural analysis by evaluating informativeness and coherence against curated standards. However, as these features derive primarily from word frequencies or local n-gram matching, they are incapable of modeling temporal topic development or the consistency.
@@ IV-C DTM-based Temporal Analysis
To understand how DTM captures these temporal dynamics, it is important to first grasp the basics of topic modeling. Topic modeling provides a probabilistic framework to uncover the latent semantic structure within textural data. While the arbitrarily generated text may be randomly distributed across documents, the topic model assumes that words in natural languages are distributed around meaningful semantic topics. Specifically, document generation is assumed to be based on a probabilistic combination of underlying topics, where each topic is represented by a probabilistic distribution over vocabulary words. In the classical Latent Dirichlet Allocation (LDA), these distributions are assumed to be static and drawn from Dirichlet priors.
Dynamic Topic Model (DTM) is an extension of classical topic models. It introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing time dependency $t$ in topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
@@
In order to apply the DTM model to narrative analysis for NCD detection, we evenly split each participant’s ASR transcription into 15 slices, assuming that each image leads to approximately similar number of words in spoken narratives. To mitigate fragmented sentences, we shift the split points to the nearest punctuation boundaries. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
@@
Figure 4: (a) The proposed model TITAN. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN, where the 15 textual chunks are manually segmented and aligned to the corresponding images. TITAN takes images and text narratives as inputs to predict NCDs, with Attention using RoPE to focus on important positions.
@@
Fig. 3 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 3(a), include “home” and “garden” in topic “A” (corresponds to settings of images 1-4), “ant”, “tree hollow” and “goose” in topic “B” (corresponds to images 5-8), “cow” and “stone” in topic “C” (corresponds to images 9-10), “fall” and “pond” in topic “D” (corresponds to images 10-11), and “many” (rabbits) and “mother” in topic “E” (corresponds to images 12-15). These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 3(b), are more chaotic. For example, “tree hollow” in topic “B” and “ant/squirrel” in topic “C” both predominantly appear in the same scene in images 5 and 6, indicating overlapping topic assignments.
@@
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC individuals mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This implies that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the story to a satisfying closure.
For the dynamic proportions of these topics over time, as shown in Fig. 3(c, d), it is observable that the HC group exhibits more distinct peaks (with each topic dominating consecutive slices) and smoother transitions between topics over time (peaks rising sharply after the preceding topic decreases), while the NCD group shows more erratic changes with less pronounced peaks (e.g., reduced peak amplitudes and sporadic dominance).
@@
Based on these findings, we proposed 6 DTM-related statistics for NCD detection. The first metric, termed “topic_consistency”, is derived from the hit rate of the top-10 topic words each time, which are derived from a DTM pretrained on all narratives. This statistic measures how consistently the subject’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). For example, if the words in text slice include all top-10 topic words at the same timestamp $t$, then topic_consistency($t$) = 1.0. Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “topic_cycle”, which is simply a boolean value indicating whether the word “home” appears in the latter half of each narrative.
The other four metrics are derived from the dynamic properties of topic evolution, including standard deviation (“topic_variability”), correlation (“topic_temporal_corr”) of the proportions over time, peak-to-peak range (difference between the highest and lowest values, termed as “topic_ptp_range”), and rate of change (averaged Euclidean distance of adjacent topic distributions over time, termed as “topic_change_rate”). These statistics capture topic development of VSNs and could be used as features followed by a downstream classifier for NCD detection, which are interpretable through their well-defined mathematical properties.
@@ IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and VSNs. This limitation impairs topic representation model, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [41]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [81, 82] (which measure the alignment between text and images through contrastive learning), we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 4.
@@
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising textual-encoder and visual-encoder to process different modalities. In this work, we adopted the visual-encoder ViT and textual-encoder RoBERTa from the pretrained Chinese-CLIP model, as mentioned in Section IV-B. We adopt this model because it can effectively align the embeddings of images and Cantonese text in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the embeddings $E_{I}$ (from the [CLS] token) and $E_{T}$ (via mean pooling) from the final projection layer of their respective encoders:
@@
where $H$ is the hidden dimension of the extracted embeddings, $J$ is the number of image frames (e.g., $J=1$ for the ADReSS dataset and $J=15$ for the CU-MARVEL-RABBIT corpus), and $K$ is the temporal length of the investigated text sequence.
@@
$E_{attn}=\text{softmax}\left(\text{Mask}(Q^{\prime}K^{\prime T})/\sqrt{H^{\prime}}\right)V$
@@
In VSN tasks such as the Rabbit Story, narrative transcriptions must align with reference images. Participants with NCDs may misidentify animal names (anomia) or miss detailed scenarios (visual perception) in these images. For example, most NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The attention module in the proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
TABLE III: Classification and Regression Results on the CU-MARVEL-RABBIT corpus. Feature dimensions are shown in parentheses. “*”: the baseline systems; $\downarrow$”: lower is better; “Emb.”: embedding(s). Best results are marked in red, and the second-best are underlined.
@@
1* Acoustics (10) SVM 0.6000 0.8054 0.4500 0.9000 0.6400 0.1580 0.3244
2* Linguistics (13) SVM 0.5376 0.6596 0.4167 0.7576 0.5700 -0.0051 0.3544
3* Micro (23) SVM 0.5843 0.7571 0.4333 0.8966 0.6300 0.1604 0.3240
4 Reference-based (16) SVM 0.6458 0.7946 0.5167 0.8611 0.6600 0.2112 0.3140
5 DTM-based (6) SVM 0.6522 0.7825 0.5000 0.9375 0.6800 0.1952 0.3172
6 Macro (22) SVM 0.6947 0.8029 0.5500 0.9429 0.7100 0.2233 0.3116
7 All statistics (45) SVM 0.6875 0.7783 0.5500 0.9167 0.7000 0.2367 0.3089
@@
where $C$ is the prediction dimension (number of classes for classification tasks and 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the fully connected layer’s weight and bias, and $\sigma$ is an activation function (softmax for classification tasks and identity for the regression task).
@@
The TITAN model, with the combination of feature extraction, embedding fusion, and Attention-based contextual alignment, can effectively integrate text and image information to evaluate the consistency and temporal flow in narratives. As shown in Fig. 4(c), the heatmap illustrates how the model captures the alignment between text chunks (x-axis) and a sequence of images (y-axis). The varying intensity of the colors represents the degree of alignment between the image and textual content, and the diagonal intensity represents the temporal consistency and flow of the narrative. Bright spots indicate high alignment, showing that text chunks and images correlate strongly at those points, while the diagonal progression reflects the sequence in which events or scenes unfold coherently over time. Moreover, compared to Fig. 3(c, d) using discrete, term-based topics (e.g., limited to 5 coarse topics), Fig. 4(c) leverages instance-level image-text matching. This contrast demonstrates that representing evolving topics with images enables a finer-grained temporal alignment between narratives and their corresponding visual concepts. These heatmaps also provide clinicians a clear view into how textual narratives and visual elements are cohesively aligned to support the model’s evaluation of narrative consistency and temporal flow.